For Exercises 1–21, find dy/dx. Assume a, b, c are constants.

x^2 + y^3 = 8

x^2y − 2y +5=0

√x + √y = 25

ax^2 − by^2 = c^2
Answer #1

x + y 2 = 8 ) o line (x² + y 3) = 0 + 2x + 3y d y = 0 - 2y ts=0 d (ay - 2y + 5) = 0 y al or (2²) + az dy - 2 dl by to 74 - C